Choose the best answer or correct letter.


1. Mary does a book with a mass of 0.5 kg from a height 50m. What is the kinetic energy if the book just before strikes the ground?

 a. 0

 b. 24.4 joules

 c. 49. joules

 d. 245 joules

2. A simple pendulum has a period of 1s. What is the length of the string?

 a. 1s

 b. 3s

 c. 2s

 d. 4s


1
Expert's answer
2021-02-28T07:37:07-0500

1)

"KE=PE=mgh=0.5\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot50\ m=245\ J."

2)

"T=2\pi\sqrt{\dfrac{L}{g}},""T=2\pi\sqrt{\dfrac{1\ m}{9.8\ \dfrac{m}{s^2}}}=2.0\ s."

Answer:

1) d. 245 joules

2) c. 2s